two trees and ash and an oak, grew together side by side in the forest. Over the years the ash grew to be one and a quarter times as high as the oak. One day the woodcutter came along and lopped 3 meters off the height of both trees and now the ash is one and a third times the height of the oak. How high was the ash originally
Answer by stanbon(75887)

An ash and an oak grew together side by side in the forest. Over the years the ash grew to be one and a quarter times as high as the oak. One day the woodcutter came along and lopped 3 meters off the height of both trees and now the ash is one and a third times the height of the oak. How high was the ash originally
-----------------
"One Day" DATA:
Ht. of oak = x ft. ; Ht of ash = (5/4)x ft
------------
"After Cut" DATA:
Ht of oak = x-3 ft ; Ht of ash = (5/4)x-3
----------------
EQUATION:
ash = (4/3)oak
(5/4)x-3 = (4/3)(x-3)
Multiply both sides by 12 to get rid of the denominators:
15x-36 = 16x-48
x=12 ft (original ht of oak)
(5/4)x = (5/4)12 = 15 ft. (original ht of ash)
============
Cheers,
Stan H.